The value of- (1+ tan11') (1+ tan34')/(1+ tan17') (1+ tan28')

Grade:12th Pass

1 Answers

divyansh sethi
12 years ago

remember if a+b=45 ,then (1+tanA)(1+tanB)=2

since,in (1+tanA)(1+tanB) ,A=11 and B =34 hence a+b=45,

similarly  2/2=1    ANS.